# Benchmarking Universal Machine Learning Interatomic Potentials for Supported Nanoparticles: Decoupling Energy Accuracy from Structural Exploration

### v1

@@ Abstract
Supported nanoparticle catalysts are widely used in the chemical industry. Computational modeling of supported nanoparticles based on density functional theory (DFT) often involves structural searches of stable local minimum energy configurations and molecular dynamics simulations at finite temperature. These are computationally demanding tasks that are intractable within DFT for large systems. In the last two decades, machine learning interatomic potentials (MLIPs) have been successfully used to substantially increase the size and time scales accessible to simulations that retain DFT accuracy. However, training reliable MLIPs is non-trivial as it requires many costly DFT calculations. Recently, several universal MLIPs (uMLIPs) have been developed, which are trained on large datasets that cover a wide range of molecules and materials. Here, we benchmark the accuracy and the efficiency of these uMLIPs in describing Cu nanoparticles supported on Al2O3 surfaces against our domain-specific DP-UniAlCu model. We find that the MACE-OMAT can reproduce reasonably well the low-energy configurations found in global optimization at an energy accuracy comparable to DP-UniAlCu. Interestingly, the MatterSim-v1.0.0-1M model, which exhibits larger deviations in the binding energies, can find even more stable configurations than the other two models in some supported nanoparticle sizes, showing its capability in structure exploration. For MD simulations, MACE-OMAT and MatterSim-v1.0.0-1M can qualitatively reproduce the mean-squared displacements of Cu atoms (MSDCu) predicted by DP-UniAlCu, albeit at roughly two orders of magnitude higher cost. We demonstrate that the uMLIPs can be very useful in simulating supported nanoparticles even without any fine-tuning, though their reduced efficiency remains a limiting factor for large-scale simulations.
@@ 1 Introduction
Atomic simulation is an indispensable tool in understanding various physical phenomena from an atomic perspective. Heterogeneous catalysis, a complex process that usually involves small molecules and solid surfaces, has benefited from valuable insights obtained from atomic simulation.Chen2020ChemRev The proper description of catalytic reactions and catalyst structures requires a reliable method at the quantum mechanical level. Density functional theory (DFT) has been the workhorse for atomic simulation over the years. However, due to its intensive computational cost, DFT-based simulations cannot satisfy the growing interest in the realistic modeling of catalytic systems that require both extended spatial and time scales.Grajciar2018ChemSocRev
@@
Supported nanoparticle catalyst is one of the most widely used catalysts in the chemical industry.Liu2018ChemRev The computational modeling of supported nanoparticles often involves the structure search of stable configurations at 0 K and the molecular dynamics (MD) simulation at finite temperature. Both the structure search and the MD are computationally demanding and become intractable to DFT when the system size increases. The typical size of a supported nanoparticle synthesized in experiments is in the range of 1-10 nm, where a reasonable structure model of a 2-nm-sized nanoparticle together with the substrate may contain thousands of atoms. The supported nanoparticle structure adopted in DFT-based studies is usually a simplified model that only contains a few dozen atoms. Thus, the computational results are difficult to directly compare with experiments. To understand and design better supported nanoparticle catalysts computationally, an accurate and efficient simulation method is highly desirable.
@@
In the last two decades, machine learning interatomic potentials (MLIPs) have been developed to accelerate atomic simulations, which retain DFT accuracy at a force-field-like cost, and find extensive applications in heterogeneous catalysis.Cheng2024PrecisChem, Omranpour2025ACSCatal, Xia2025ChemSocRev However, the development of an MLIP is a non-trivial task, which requires the user to collect a large amount of ab initio data through many training-validation-application cycles. Also, the MLIP is usually specific to one system, and its transferability is limited. For example, an MLIP trained on one crystalline phase may be useful for exploring configurations of another phase, but the energy and the forces are normally less accurate. Recently, several universal MLIPs (uMLIPs) have emerged, namely, DPA2Zhang2024npjComputMater, DPA3,Zhang2025arXiv MACE-MPBatatia2025JCP, and MatterSimYang2024arXiv_mattersim. These uMLIPs are trained on large datasets that cover a wide range of molecules and materials, such as Materials Project (MP),Jain2013APLMater Open Catalyst Project (OC2M),Tran2023ACSCatal and Open Materials Dataset (OMAT),BarrosoLuque2024arXiv and are expected to apply to a wide range of systems. In general, uMLIPs provide a good foundation to develop an MLIP for a specific system since they can generate reasonable configurations without active learning and can be further fine-tuned to a target accuracy if necessary. MACE-MPBatatia2025JCP has been tested on numerous systems from solid materials to heterogeneous interfaces, while other recent uMLIPs have been mainly tested with solid-state systems. Despite all the tests that have been done, the benchmark of the uMLIPs on supported nanoparticles is still lacking.
@@
In this work, we examined several uMLIPs on simulating supported nanoparticles, where two main application scenarios are demonstrated: (1) global optimization that produces low-energy configurations of nanoparticles serving good initial structures for the further investigation of dynamics, and (2) MD that reveals the dynamics of supported nanoparticles at finite temperature. We choose Cu nanoparticles on Al2O3 surfaces as the benchmark system and set the domain-specific deep potential (DP) model developed in our previous workXu2025ACSNano as the benchmark baseline, which is denoted as DP-UniAlCu in the following sections. The DP-UniAlCu model is trained on a dataset named UniAlCu, including various structures of supported Cu nanoparticles on Al2O3 surfaces. The uMLIPs examined by this work include DPA2 (mptrj and oc20),Zhang2024npjComputMater DPA3 (omat24),Zhang2025arXiv MACE-MP (medium),Batatia2025JCP MACE-OMAT (medium),BarrosoLuque2024arXiv and MatterSim (v1.0.0-1M and v1.0.0-5M).Yang2024arXiv_mattersim
@@ 2 Results
Minimum Energy Configurations from Cu1 to Cu21. We commence our examination with geometric optimization-based tasks. Figure 1 shows the binding energy (defined in Equation 1) of Cu1-21 on three Al2O3 surfaces by various MLIPs, which are obtained from global optimization in our previous work.Xu2025ACSNano The three Al2O3 surfaces are modeled by $p(3\times 2)$ $\gamma$-Al2O3(100), $p(2\times 2)$ $\gamma$-Al2O3(110), $p(3\times 2\sqrt{3})$ $\alpha$-Al2O3(0001) periodic slabs (Figure S1). The comparison of the absolute total energy is not accessible as these MLIPs are trained on datasets labelled by different DFT calculation settings. The Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional is used for the reported DFT binding energy results and also the dataset to train our DP-UniAlCu model. MACE and MatterSim-branched uMLIPs are trained on datasets labelled by DFT-PBE using a similar setting as ours (Table S1 and S2), which is the default one named as “MPRelaxSet” used in Materials Project.Jain2013APLMater, Ong2013ComputMaterSci The binding energy results from our DFT setting and “MPRelaxSet” is almost identical with a maximum absolute error of 0.10 eV (Figure S2). Since all MLIPs tested here are trained on DFT-PBE, we are able to examine the accuracy of these MLIPs on the binding energy, where a reliable MLIP should be able to reproduce the binding energy trend across different nanoparticle sizes and surfaces. We noticed the dispersion interaction may affect the absolute values of binding energies of supported nanoparticles when compared to experiments,Wellendorff2015SurfSci, Zhao2024ACSCatal however, the DFT-PBE results presented here only served as a reference for the comparison of different MLIPs and the relative binding strength across different nanoparticle sizes and supporting surfaces is preserved.Xu2025ACSNano Figure 1a-c show the binding energy results of Cu1-21 on $\gamma$-Al2O3(100), $\gamma$-Al2O3(110), and $\alpha$-Al2O3(0001), respectively. The MLIP binding energy results are calculated based on the MLIP total energies of the configurations minimized by the corresponding MLIP model starting from the DFT-minimized configurations. No significant structural difference between DFT and MLIP-minimized configurations is observed. Most models give minimized structures with an averaged atomic displacement smaller than $\sim$0.05 Å compared to the DFT-minimized structures except for DPA2-MPTrj and DPA2-OC2M that give displacements larger than $\sim$0.1 Å (Figure S3). All examined uMLIPs can reproduce the binding energy trend as a function of the nanoparticle size on a single Al2O3 surface, namely, the binding energy on $\gamma$-Al2O3(100) is almost independent of the nanoparticle size while the binding energies on $\gamma$-Al2O3(110) and $\alpha$-Al2O3(0001) decrease significantly with the nanoparticle size. When comparing the binding energies of the same nanoparticle size across different Al2O3 surfaces, some uMLIPs cannot capture the relative binding strength to the supporting surface, where the trend by DFT-PBE from the strongest to the weakest should be $\gamma$-Al2O3(110), $\gamma$-Al2O3(100), and $\alpha$-Al2O3(0001). For example, DPA2-MPTrj predicts a stronger binding on $\alpha$-Al2O3(0001) than on $\gamma$-Al2O3(110), which is inconsistent with DFT-PBE. Figure 1d shows the distribution of binding energy errors through the violin plots, where a good model should have an error distribution centered at zero with a small spread. The binding energy error is defined as the difference between the binding energy predicted by an MLIP and the binding energy by DFT-PBE. MACE-OMAT and MatterSim-v1.0.0-5M exhibit small errors with narrow spreads centered near zero, which reproduce both the relationship between the binding energy and the nanoparticle size and the relative binding strength across three Al2O3 surfaces. It is surprising that MACE-OMAT and MatterSim-v1.0.0-5M are not trained explicitly on any structures of supported nanoparticles but have an accuracy comparable to our DP-UniAlCu model. MACE-MP, which has the same architecture as MACE-OMAT but is trained on a smaller dataset, is less accurate than MACE-OMAT and predicts more negative binding energies on all three Al2O3 surfaces, which can be seen from the average binding energy errors. MatterSim-v1.0.0-1M, which is trained on the same dataset as MatterSim-v1.0.0-5M but has a smaller number of model parameters, leads to a degraded accuracy on binding energy, particularly for $\gamma$-Al2O3(110) with the averaged binding energy error of $\sim$2 eV. The above results indicate that the model performance can be improved either by increasing the dataset size (MACE-MP vs. MACE-OMAT) or by increasing the model size (MatterSim-v1.0.0-1M vs. MatterSim-v1.0.0-5M). To examine this model improvement strategy, we trained a MACE model based on our UniAlCu dataset using the same architecture as MACE-OMAT, which is denoted as the MACE-UniAlCu. The details of the training settings and the model errors on the entire UniAlCu dataset are given in Section S3. MACE-UniAlCu exhibits smaller error spreads for all three Al2O3 surfaces compared to both DP-UniAlCu and MACE-OMAT. The DPA2-MPTrj shows much larger errors compared to other MLIP models. The large errors are also observed for the DPA2-OC2M (Figure S5), which is trained on the OC2M dataset.Chanussot2021ACSCatal The DPA3-OMAT exhibits smaller errors than DPA2-MPTrj and DPA2-OC2M due to its improved architecture. The model architecture can affect the accuracy, as seen from that DPA2-MPTrj is the head trained on the same dataset as MACE-MP, but has larger errors, which is similar in the case of DPA3-OMAT and MACE-OMAT. In addition, the multi-head training strategy used in DPA2 and DPA3, despite utilizing available datasets to the maximum extent, compromises the accuracy of a specific head due to inconsistent DFT settings in the datasets.
@@
Global Optimization of Larger Nanoparticles. In practice, the global minimum configurations of supported nanoparticles are not available and are usually difficult to construct manually. Global optimization can explore low-energy configurations and possibly find the global minimum configurations of supported nanoparticles, where the initial structures are randomly generated, and the final structures can have different geometries but similar energies. A good MLIP should be able to minimize the initial random structures that may have unphysically short interatomic distances and distinguish the low-energy configurations from the high-energy ones. Therefore, global optimization is a good test for the uMLIPs to examine their stability in minimization and their performance in finding low-energy configurations. Based on the performance of uMLIPs on the binding energies of Cu1-21, we examined MACE-OMAT and MatterSim-V1.0.0-1M due to their balance in accuracy and efficiency. At the same time, we tested DP-UniAlCu as a baseline. The global optimization based on the genetic algorithm was performed for several larger nanoparticles, including Cu27, Cu38, Cu47 and Cu55. All the combinations of nanoparticle sizes and Al2O3 surfaces are not included in our UniAlCu dataset, making these global optimization tasks good examinations for the model’s extrapolation ability. The periodic slabs of $p(5\times 3)$ $\gamma$-Al2O3(100), $p(3\times 3)$ $\gamma$-Al2O3(110), $p(5\times 3\sqrt{3})$ $\alpha$-Al2O3(0001) shown in Figure S6 are used to accommodate the nanoparticles. In the global optimization task, the initial structures are randomly generated, and the structures in the following 20 generations are generated based on cut-and-splice and mutation algorithms, where the population size is set to 50. The details of the genetic algorithm are given in Section Methods. Among all obtained MLIP-minimized configurations, 50 structures with the lowest energies are further optimized by DFT. For the three MLIPs examined on one nanoparticle-surface combination, there will be 150 DFT-minimized configurations in total, forming the low-energy ensemble.
Table 1 gives the error distributions on the low-energy ensembles. For nanoparticles on the two $\gamma$-Al2O3 surfaces, our DP-UniAlCu model and the MACE-OMAT model have smaller error spreads compared to MatterSim-v1.0.0-1M. Our DP-UniAlCu is slightly better as the MACE-OMAT model gives more negative binding energies and exhibits larger error spreads, for example, Cu38 on $\gamma$-Al2O3(100). The Mattersim-v1.0.0-1M model gives significantly large errors for Cu27, Cu38, and Cu47 on two $\gamma$-Al2O3 surfaces and predicts decreasing trends in the binding energy, while it gives results of Cu55 closer to DFT. This non-uniform accuracy indicates that the extrapolation ability of a uMLIP can be strongly system-dependent and cannot be guaranteed for consistency across systems, which may be rooted deeply in their training datasets. The DP-UniAlCu and MACE-OMAT models exhibit similar non-uniform accuracies but less severe, for example, they are slightly less accurate for two $\gamma$-Al2O3 surfaces than $\alpha$-Al2O3(0001). Though large offsets are observed, the error spreads by MatterSim-v1.0.0-1M are generally small, indicating the relative stability of the low-energy configurations may pertain. For $\alpha$-Al2O3(0001), all three models exhibit larger error spreads. The error averages by DP-UniAlCu and MACE-OMAT are close to zero, while the error average by MatterSim-v1.0.0-1M is around +2 eV. For Cu47 and Cu55, MACE-OMAT performs better than DP-UniAlCu, showing smaller error spreads.
[…]

### v2

@@ Abstract
Supported nanoparticle catalysts are widely used in the chemical industry. Computational modeling of supported nanoparticles based on density functional theory (DFT) often involves structural searches of stable local minimum energy configurations and molecular dynamics simulations at finite temperature. These are computationally demanding tasks that are intractable within DFT for large systems. In the last two decades, machine learning interatomic potentials (MLIPs) have been successfully used to substantially increase the size and time scales accessible to simulations approximating DFT accuracy. However, training reliable MLIPs is non-trivial as it requires many costly DFT calculations. Recently, several universal MLIPs (uMLIPs) have been developed, which are trained on large datasets that cover a wide range of molecules and materials. Here, we benchmark the accuracy and the efficiency of these uMLIPs in describing Cu nanoparticles supported on Al2O3 surfaces against our domain-specific DP-UniAlCu model. We find that the MACE-OMAT can reproduce reasonably well the low-energy structures found in global optimization at an energy accuracy comparable to DP-UniAlCu. Interestingly, the MatterSim-v1.0.0-1M model, which exhibits larger deviations in the binding energies, can find even more stable configurations than the other two models in some supported nanoparticle sizes, showing its capability in structure exploration. For MD simulations, MACE-OMAT and MatterSim-v1.0.0-1M can qualitatively reproduce the mean-squared displacements of Cu atoms (MSDCu) predicted by DP-UniAlCu, albeit at roughly two orders of magnitude higher cost. We demonstrate that the uMLIPs can be very useful in simulating supported nanoparticles even without any fine-tuning, though their reduced efficiency remains a limiting factor for large-scale simulations.
@@ 1 Introduction
Atomic simulation is an indispensable tool in understanding various physical phenomena from an atomic perspective. Heterogeneous catalysis, a complex process that usually involves small molecules and solid surfaces, has benefited from valuable insights obtained from atomic simulation.5 The proper description of catalytic reactions and catalyst structures requires a reliable method at the quantum mechanical level. Density functional theory (DFT) has been the workhorse for atomic simulation over the years. However, due to its intensive computational cost, DFT-based simulations cannot satisfy the growing interest in the realistic modeling of catalytic systems that require both extended spatial and time scales.10
@@
Supported nanoparticle catalyst is one of the most widely used catalysts in the chemical industry.18 The computational modeling of supported nanoparticles often involves the structure search of stable configurations at 0 K and the molecular dynamics (MD) simulation at finite temperature. Both the structure search and the MD are computationally demanding and become intractable to DFT when the system size increases. The typical size of a supported nanoparticle synthesized in experiments is in the range of 1-10 nm,7 where a reasonable structure model of a 2-nm-sized nanoparticle together with the substrate may contain thousands of atoms. The supported nanoparticle structure adopted in DFT-based studies is usually a simplified model that only contains a few dozen atoms. Thus, the computational results are difficult to directly compare with experiments. To understand and design better supported nanoparticle catalysts computationally, an accurate and efficient simulation method is highly desirable.
@@
In the last two decades, machine learning interatomic potentials (MLIPs) have been developed to accelerate atomic simulations, which approximate DFT accuracy at almost force-field cost, and find extensive applications in heterogeneous catalysis.6, 20, 28 However, the development of an MLIP is a non-trivial task, which requires the user to collect a large amount of ab initio data through many training-validation-application cycles. Also, the MLIP is usually specific to one system, and its transferability is limited. For example, an MLIP trained on one crystalline phase may be useful for exploring configurations of another phase, but the energy and the forces are normally less accurate. Recently, several universal MLIPs (uMLIPs) have emerged, namely, DPA233, DPA3,34 MACE-MP2, and MatterSim32. These uMLIPs are trained on large datasets that cover a wide range of molecules and materials, such as Materials Project (MP),11 Open Catalyst Project (OC2M),24 and Open Materials Dataset (OMAT),1 and are expected to apply to a wide range of systems. In general, uMLIPs provide a good foundation to develop system specific MLIPs because they can generate meaningful configurations without active learning and can be further fine-tuned to a target accuracy if necessary.19 MACE-MP2 has been tested on numerous systems from solid materials to heterogeneous interfaces, while other recent uMLIPs have been mainly tested with solid-state systems. Despite all the tests that have been done, the benchmark of the uMLIPs on supported nanoparticles is still lacking.
@@
In this work, we examined several uMLIPs on simulating supported nanoparticles, where two main application scenarios are demonstrated: (1) global optimization that produces low-energy structures of nanoparticles serving good initial structures for the further investigation of dynamics, and (2) MD that reveals the dynamics of supported nanoparticles at finite temperature. We choose Cu nanoparticles on Al2O3 surfaces as the benchmark system and set the domain-specific deep potential (DP) model developed in our previous work29 as the benchmark baseline, which is denoted as DP-UniAlCu in the following sections. The dataset used to train DP-UniAlCu comprises 147,464 structures in total, both at zero and at finite temperatures ranging from 400 K to 1200 K, and includes Cu structures (bulk, surface, and gas-phase nanoparticle), Al2O3 surfaces, and Al2O3-supported Cu nanoparticles. The dataset is publicly available in a zenodo repository.30 The uMLIPs examined by this work include DPA2 (mptrj and oc20),33 DPA3 (omat24),34 MACE-MP (medium),2 MACE-OMAT (medium),1 and MatterSim (v1.0.0-1M and v1.0.0-5M).32
@@ 2 Results
Minimum Energy Configurations from Cu1 to Cu21. We commence our examination with geometric optimization-based tasks. The binding energies (defined in Equation 1) of Cu1-21 supported on three Al2O3 surfaces are evaluated by various MLIPs. The corresponding structures were obtained from global optimization in our previous work.29 The three Al2O3 surfaces are modeled by $p(3\times 2)$ $\gamma$-Al2O3(100), $p(2\times 2)$ $\gamma$-Al2O3(110), $p(3\times 2\sqrt{3})$ $\alpha$-Al2O3(0001) periodic slabs (Figure S1). The comparison of the absolute total energy is not accessible as these MLIPs are trained on datasets labelled by different DFT calculation settings. The Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional is used for the reported DFT binding energy results and also the dataset to train our DP-UniAlCu model. MACE and MatterSim-branched uMLIPs are trained on datasets labelled by DFT-PBE using a similar setting as ours (Table S1 and S2), which is the default one named as “MPRelaxSet” used in Materials Project.A. Jain, S. P. Ong, G. Hautier, W. Chen, W. D. Richards, S. Dacek, S. Cholia, D. Gunter, D. Skinner, G. Ceder, et al. (2013), 22 The binding energy results from our DFT setting and “MPRelaxSet” is almost identical with a maximum absolute error of 0.10 eV (Figure S2). Since all MLIPs tested here are trained on DFT-PBE, we are able to examine the accuracy of these MLIPs on the binding energy, where a reliable MLIP should be able to reproduce the binding energy trend across different nanoparticle sizes and surfaces. We noticed the dispersion interaction may affect the absolute values of binding energies of supported nanoparticles when compared to experiments,27, 36 however, the DFT-PBE results presented here only served as a reference for the comparison of different MLIPs and the relative binding strength across different nanoparticle sizes and supporting surfaces is preserved.29 Figure 1a-c show the binding energy results of Cu1-21 on $\gamma$-Al2O3(100), $\gamma$-Al2O3(110), and $\alpha$-Al2O3(0001), respectively. The MLIP binding energy results are calculated based on the MLIP total energies of the configurations minimized by the corresponding MLIP model starting from the DFT-minimized configurations. No significant structural difference between DFT and MLIP-minimized configurations is observed. Most models give minimized structures with an averaged atomic displacement smaller than $\sim$0.05 Å compared to the DFT-minimized structures except for DPA2-MPTrj and DPA2-OC2M that give displacements larger than $\sim$0.1 Å (Figure S3). All examined uMLIPs can reproduce the binding energy trend as a function of the nanoparticle size on a single Al2O3 surface, namely, the binding energy on $\gamma$-Al2O3(100) is almost independent of the nanoparticle size while the binding energies on $\gamma$-Al2O3(110) and $\alpha$-Al2O3(0001) decrease significantly with the nanoparticle size. The size dependence of the binding energy of the nanoparticles originates from the distinct surface sites present on the three Al2O3 surfaces. In particular, the presence of strong AlIII binding sites on $\gamma$-Al2O3(110) and $\alpha$-Al2O3(0001) (Figure S1) enables an increasing number of Cu atoms to interact strongly with the support as the nanoparticle size grows, resulting in progressively stronger binding, as discussed in our previous work.29 According to DFT-PBE, the binding strength follows the trend $\gamma$-Al2O3(110) >> $\alpha$-Al2O3(0001) >> $\gamma$-Al2O3(100). The three panels of Figure S4 show the difference in the binding energy of the $N$-atom nanoparticles between pairs of surfaces. All uMLIPs reproduce the sign of the binding energy difference, with the exception of DPA2-MPTrj and DPA2-OC2M. For example, DPA2-MPTrj predicts stronger binding on $\alpha$-Al2O3(0001) than on $\gamma$-Al2O3(110), contrary to the DFT-PBE reference. Figure 1d shows the distribution of binding energy errors through the violin plots, where a good model should have an error distribution centered at zero with a small spread. The binding energy error is defined as the difference between the binding energy predicted by an MLIP and the binding energy by DFT-PBE. MACE-OMAT and MatterSim-v1.0.0-5M exhibit small errors with narrow spreads centered near zero, which reproduce both the relationship between the binding energy and the nanoparticle size and the relative binding strength across three Al2O3 surfaces. It is surprising that MACE-OMAT and MatterSim-v1.0.0-5M are not trained explicitly on any structures of supported nanoparticles but have an accuracy comparable to our DP-UniAlCu model. MACE-MP, which has the same architecture as MACE-OMAT but is trained on a smaller dataset, is less accurate than MACE-OMAT and predicts more negative binding energies on all three Al2O3 surfaces, which can be seen from the average binding energy errors. MatterSim-v1.0.0-1M, which is trained on the same dataset as MatterSim-v1.0.0-5M but has a smaller number of model parameters, leads to a degraded accuracy on binding energy, particularly for $\gamma$-Al2O3(110) with the averaged binding energy error of $\sim$2 eV. The above results indicate that the model performance can be improved either by increasing the dataset size (MACE-MP vs. MACE-OMAT) or by increasing the model size (MatterSim-v1.0.0-1M vs. MatterSim-v1.0.0-5M). To examine this model improvement strategy, we trained a MACE model based on our UniAlCu dataset using the same architecture as MACE-OMAT, which is denoted as the MACE-UniAlCu. The details of the training settings are given in Section S3 and the model errors on the entire UniAlCu dataset are shown in Figure S5. MACE-UniAlCu exhibits smaller error spreads for all three Al2O3 surfaces compared to both DP-UniAlCu and MACE-OMAT. The DPA2-MPTrj shows much larger errors compared to other MLIP models. The large errors are also observed for the DPA2-OC2M (Figure S6), which is trained on the OC2M dataset.4 The DPA3-OMAT exhibits smaller errors than DPA2-MPTrj and DPA2-OC2M due to its improved architecture. The model architecture can affect the accuracy, as seen from that DPA2-MPTrj is the head trained on the same dataset as MACE-MP, but has larger errors, which is similar in the case of DPA3-OMAT and MACE-OMAT. In addition, the multi-head training strategy used in DPA2 and DPA3, despite utilizing available datasets to the maximum extent, compromises the accuracy of a specific head due to inconsistent DFT settings in the datasets.
@@
Global Optimization of Larger Nanoparticles. In practice, the low-energy structures of supported nanoparticles are not available and are usually difficult to construct manually. Global optimization can explore low-energy structures of supported nanoparticles, where the initial structures are randomly generated, and the final structures can have different geometries but similar energies. A good MLIP should be able to minimize the initial random structures that may have unphysically short interatomic distances and distinguish the low-energy structures from the high-energy ones. Therefore, global optimization is a good test for the uMLIPs to examine their stability in minimization and their performance in finding low-energy structures. Based on the performance of uMLIPs on the binding energies of Cu1-21, we examined MACE-OMAT and MatterSim-v1.0.0-1M due to their balance in accuracy and efficiency. At the same time, we tested DP-UniAlCu as a baseline. The global optimization based on the genetic algorithm was performed for several larger nanoparticles, including Cu27, Cu38, Cu47 and Cu55. Our UniAlCu training dataset only included supported nanoparticles up to Cu21. Thus, global optimization of larger supported nanoparticles, such as Cu27, Cu38, Cu47, and Cu55, and comparison of their energy with that of the corresponding DFT structures is a test of the extrapolation ability of the model. The periodic slabs of $p(5\times 3)$ $\gamma$-Al2O3(100), $p(3\times 3)$ $\gamma$-Al2O3(110), $p(5\times 3\sqrt{3})$ $\alpha$-Al2O3(0001) shown in Figure S7 are used to accommodate the nanoparticles. In the global optimization task, the initial structures are randomly generated, and the structures in the following 20 generations are generated based on cut-and-splice and mutation algorithms, where the population size is set to 50. The details of the genetic algorithm are given in Section Methods. Since the search is finite, the minimum energy structures are unlikely to be the true global minima. Let $E_{b}$ be the binding energies of the MLIP minimum energy structures. After local DFT relaxation the structures differ little from their MLIP counterparts and have energies $E_{b}^{\mathrm{DFT}}$. For each nanoparticle-surface combination this yields 150 MLIP low energy structures in one to one correspondence with 150 structures locally optimized by DFT. We define the binding-energy deviation by $\Delta E_{b}=E_{b}-E_{b}^{\mathrm{DFT}}$.
Table 1 summarizes the energetic metrics of the low-energy structures. For each model, we report the average binding energy deviation, $\mathrm{avg}(\Delta E_{b})$, which should approach zero when the model exactly reproduces DFT, and its standard deviation, $\mathrm{std}(\Delta E_{b})$. These two quantities measure the bias and the spread introduced by the models. For nanoparticles supported on the two $\gamma$-Al2O3 surfaces, MACE-OMAT yields $\Delta E_{b}$ values closer to zero among the three models. These results suggest that MACE-OMAT maintains competitive predictive performance for larger supported nanoparticles, even though it was not explicitly trained on supported nanoparticle structures. In comparison, MatterSim-v1.0.0-1M has $\Delta E_{b}$ values comparable to DP-UniAlCu for $\gamma$-Al2O3(100), but makes substantially larger errors for $\gamma$-Al2O3(110). For $\alpha$-Al2O3(0001), DP-UniAlCu provides the best accuracy, while MACE-OMAT and MatterSim-v1.0.0-1M exhibit larger deviations. Overall, each model achieves its best performance on a different support: DP-UniAlCu on $\alpha$-Al2O3(0001), MACE-OMAT on $\gamma$-Al2O3(110), and MatterSim-v1.0.0-1M on $\gamma$-Al2O3(100). This non-uniform accuracy indicates that the extrapolation performance of the MLIPs is not consistent across different supports, likely reflecting differences in training data coverage and protocols. To further examine this behavior, we locally relaxed the above 150 structures with MACE-UniAlCu to obtain the corresponding $\Delta E_{b}$. The energetic metrics obtained in this way are reported in Table S3, which shows that MACE-UniAlCu exhibits improved accuracy for $\gamma$-Al2O3(100) but reduced accuracy for the other two surfaces relative to DP-UniAlCu, reinforcing the conclusion that extrapolation performance varies across systems even for models trained on the same dataset.
In practice, MLIP-generated low-energy structures can be further minimized by DFT to identify the DFT low-energy structures. Therefore, the ability of an MLIP not only to include but also to preserve the relative energetic ordering of these DFT low-energy structures serves as a key indicator of its performance. Two metrics (Table 2) are used to quantify the quality of MLIP-generated low-energy structures, namely, rank biased overlap26 (RBO, defined in Equation 2) and Kendall’s $\tau$12 (defined in Equation 3). The RBO measures the similarity between two lists of binding energies, the DFT ones and the MLIP ones, and focuses more on the top of the lists, where a larger RBO value indicates a better MLIP performance. However, it is less sensitive to the order of the binding energies in the lists. In addition, Kendall’s $\tau$ measures the ordinal association between two lists of binding energies, where a larger value indicates a better MLIP performance. In seven out of twelve systems, our DP-UniAlCu model shows the best performance in both RBO and Kendall’s $\tau$, indicating its excellent ability to generate low-energy structures consistent with DFT. For the rest of the systems, MACE-OMAT performs better than DP-UniAlCu. MatterSim-v1.0.0-1M gives less satisfactory results and only exhibits comparable performance on a few systems, such as Cu47 on $\gamma$-Al2O3(100) and Cu55 on $\gamma$-Al2O3(110).
To better visualize the ability of the MLIP models to identify low-energy structures across nanoparticle sizes, for each system (nanoparticle-surface) we rank all candidate structures by their binding-energy deviation relative to the minimum energy structure among the 150 DFT-optimized structures, i.e., $\Delta\tilde{E}_{b}=E_{b}-\min(E_{b}^{\mathrm{DFT}})$. The $\Delta\tilde{E}_{b}$ spectra of the nanoparticles on the three Al2O3 surfaces are shown in Figure 2a-c for Cu27 and Cu38 and in Figure 3a-c for Cu47 and Cu55. For each system, the three spectra at the bottom of a panel show the $\Delta\tilde{E}_{b}$ values predicted by DP-UniAlCu, MACE-OMAT, and MatterSim-v1.0.0-1M, respectively, while the spectrum at the top of the panel reports the corresponding binding energies after local DFT optimization. The structures are color-coded according to the MLIP used for their generation, that is blue for DP-UniAlCu, green for MACE-OMAT, and orange for MatterSim-v1.0.0-1M. In each panel the minimum DFT binding energy ( $\min(E_{b}^{\mathrm{DFT}})$) is aligned with the zero of the energy scale and is indicated by a vertical arrow. The arrows in the MLIP spectra indicate the MLIP energy of the structure corresponding to the minimum DFT structure.
DP-UniAlCu finds the structure of minimum $E_{b}$ in eight cases, MatterSim-v1.0.0-1M in three cases, and MACE-OMAT in one case. While the DP-UniAlCu and MACE-OMAT spectra overlap significantly with each other and with the DFT-minimized spectrum, MatterSim-v1.0.0-1M is systematically shifted to higher $E_{b}$ values. Further analysis of the binding-energy error decomposition (Table S4) shows that the dominant binding-energy errors in MatterSim-v1.0.0-1M originate from the gas-phase nanoparticle energies. Since optimization of supported nanoparticles is performed at fixed composition, these errors introduce a constant offset and therefore do not affect the relative energy ranking among low-energy structures. In the four cases in which DP-UniAlCu fails to locate the structure with the lowest DFT energy, the corresponding $E_{b}$ is very close to $\min E_{b}^{\mathrm{DFT}}$. Interestingly, the DP energy of two minimum energy structures identified by MatterSim-v1.0.0-1M is also the minimum energy of the DP spectrum, indicating that the correct identification of low energy structures depends not only on the accuracy of the MLIP but also reflects the finite character and stochastic nature of a single search by the genetic algorithm. Although MatterSim-v1.0.0-1M exhibits a systematic binding energy shift, the minimum energy structures that it identifies have good ranking metrics (RBO and Kendall’s $\tau$), similar to those of DP-UniAlCu, indicating that these structures are not a random outcome of a search. This suggests that there is no strong correlation between energetic predictions and ranking metrics. To further test this hypothesis we report in Table S5 the ranking metrics of MACE-UniAlCu on the same low-energy structures. While the energy accuracy of MACE-UniAlCu is sometimes better and sometimes worse than that of DP-UniAlCu (see Table S3), MACE-UniAlCu achieves consistently improved ranking metrics for almost all the systems. These results suggest that even in the presence of significant absolute binding-energy errors a model can still predict a meaningful ordering of the low energy structures. To improve our confidence on the predicted structures it is useful to complement energetic predictions with ranking metrics.
Finite-Temperature Dynamics. The dynamic behavior of supported nanoparticles is important for understanding the catalytic activity and stability. MD trajectories are driven by forces, and we report in Table 3 the root mean-squared errors (RMSEs) of the MLIP forces relative to their DFT counterpart, using the UniAlCu dataset, which includes three subsystems, i.e., pure Cu structures, $\gamma$-Al2O3 surfaces with and without supported nanoparticles, and $\alpha$-Al2O3 surfaces with and without supported nanoparticles. Unsurprisingly, DP-UniAlCu and MACE-UniAlCu give the best force predictions. MACE-OMAT, DPA3-OMAT and MatterSim-v1.0.0-1M show similar accuracy. Several MD-based tasks are designed to examine MLIPs, as discussed below.